OpenStudy (anonymous):

convert 600 degrees to radians

OpenStudy (johnweldon1993):

To convert a measurement of degrees to radians...you multiply by pi/180 \[\large 600^{\circ} \times \frac{\pi}{180^{\circ}} = ?\]